OpenStudy (anonymous):

Indefinite Integral Help int (x/(2x+1) dx I keep getting (x/2)-((1/4)ln|2x+1|)+1/4+C Book says ((1/2)x)-((1/4)ln|2x+1|)+C. So when I get a constant as part of the answer when trying to find the indefinite integral, do I just consider it as part of the constant C? Is that why I'm off by +1/4?

OpenStudy (turingtest):

yes, 1/4 is absorbed into C, because C is unknown that will happen a lot in differential equations, so get used to it

OpenStudy (akshay_budhkar):

yes whenever you have another constant, you ought to add it into C. The C absorbs it.. If you are having marks for steps then this step will count for marks as well ( maybe 1/2 mark or so if u r scoring full elsewhere)

OpenStudy (turingtest):

wolfram keeps the 1 though ironically http://www.wolframalpha.com/input/?i=int%20(x%2F(2x%2B1)%20dx&t=crmtb01 but they left it in the parentheses, so I guess they feel it's justified

OpenStudy (akshay_budhkar):

but in the alternate forms it does absorb the constant :D

OpenStudy (turingtest):

right^

OpenStudy (akshay_budhkar):

turing i need u 1 min

OpenStudy (akshay_budhkar):

i will give u the link

OpenStudy (turingtest):

I don't even see why they add the 1 for 'restricted values of x' in wolf what does that mean, do you know akshay?

myininaya (myininaya):

what are you guys doing?

myininaya (myininaya):

comparing your answers with wolfram?

myininaya (myininaya):

sprusty 1/4+constant is still a constant so your answer and the book answer are fine! :)

OpenStudy (turingtest):

yes, but why does wolf add the 1 for 'restricted values' I want to know

myininaya (myininaya):

\[\int\limits_{}^{}\frac{x}{2x+1} dx\] u=2x+1 => du=2 dx u=2x+1 => (u-1)/2=x \[\int\limits_{}^{}x \cdot \frac{1}{2x+1} dx=\int\limits_{}^{}\frac{u-1}{2} \cdot \frac{1}{u} \frac{du}{2}\] \[\frac{1}{4}\int\limits_{}^{}\frac{u-1}{u} du=\frac{1}{4}\int\limits_{}^{}(1-\frac{1}{u}) du=\frac{1}{4}(u-\ln|u|)+C\]

myininaya (myininaya):

\[=\frac{1}{4}(2x+1-\ln|2x+1|)+C\]

myininaya (myininaya):

\[=\frac{x}{2}+\frac{1}{4}-\frac{1}{4} \ln|2x+1| +C\]

myininaya (myininaya):

\[=\frac{x}{2}-\frac{1}{4}\ln|2x+1|+c \]

myininaya (myininaya):

i will look at wolfram's solution one sec

myininaya (myininaya):

No I don't know what they are talking about honestly lol

myininaya (myininaya):

either one of those things they have listed is fine though

myininaya (myininaya):

imperialist seems smart I will ask him what he think wolfram means